OpenStudy (anonymous):

what numbers do u multiply to get -30 but add to get 12 ???

Parth (parthkohli):

\( \color{Black}{\Rightarrow ab = -30 }\) \( \color{Black}{\Rightarrow a + b = 12 }\) I think you are doing factoring

OpenStudy (anonymous):

yes

Parth (parthkohli):

I guess trial and error is better

Parth (parthkohli):

You can solve this by substitution though

OpenStudy (anonymous):

so what are the numbers?

Parth (parthkohli):

Let's go with trial and error here. And logic. One number has to be negative on the list because two negatives make a positive and two positives make a positive too.

OpenStudy (anonymous):

right...

Parth (parthkohli):

Also the number with the positive is much higher than the number with the negative.

OpenStudy (anonymous):

okay..

Parth (parthkohli):

-5 * 6 = 30 -3 * 10 = 30 -2 * 15 = 30

Parth (parthkohli):

What's the expression btw?

OpenStudy (anonymous):

yeah but none of those numbers add up to get 12

Parth (parthkohli):

Can you tell me what trinomial is to be factored?

OpenStudy (anonymous):

2x^2+12x-15

Parth (parthkohli):

I think that the only way here is to complete the square.

OpenStudy (anonymous):

okay thank you!

Parth (parthkohli):

Yw!